please help me with this question:Y/2=12

Answer by stanbon(75887)

Y/2=12
----
That equation is saying "half of y is 12".
----
To solve the equation for "y", multiply both sides of
the equation by 2.
---
2(y/2) = 2*12
Cancel the common factor of "2" on the left side to get:
y = 24
=================
Cheers,
Stan H.